OpenStudy (amenah8):

Graph first function using its second derivative f(x)=2x^3 - x^2 + 2x - 1 CLICK TO SEE MY WORK SO FAR!

OpenStudy (amenah8):

First D: 6x^2 - 2x +2 =0 --> USE THIS TO FIND MAXIMUM AND MINIMUM (how do I do that?) Second D: 12x - 2 =0 --> USE THIS TO FIND INFLECTION POINTS (how do I do that??) Next, I know I have to find the x and y intercepts, and then set D1 as <0 and >0 and the same with D2 Then I use all of that information to graph the original function

OpenStudy (danjs):

most th etime in calc , even if you forget, take the derivatives for at least some credit maybe..

OpenStudy (amenah8):

i got the first 2 derivatives. and i know the steps i have to take, i just don't know how to do that with this problem...

OpenStudy (danjs):

the derivatives look correct, just a use of power rule thing f(x)=2x^3 - x^2 + 2x - 1 f ' (x) = 6x^2 - 2x + 2 f '' (x) = 12x - 2

OpenStudy (danjs):

do you understand the first derivative is the instantaneous rate of change of the function at any point, the slope of a tangent to the curve at any point

OpenStudy (amenah8):

yes, i understand. but in class we said we next had to set each derivative (the 1st and 2nd equal to 0). somehow, with D1, we can find the max and min. And with D2, we can find the inflection point

OpenStudy (danjs):

right, the first derivative can tell you if the function is increasing or decreasing at any point on the curve. so if you set that to zero f ' (x) = 6x^2 - 2x + 2 = 0 the slope at those x values will be 0. a horizontal line.. which occurs at the apex or a man value or a min value most the time, you have to test if the function changes from increasing to decreasing at those x values...

OpenStudy (danjs):

you see?

OpenStudy (amenah8):

to test it, do i have to factor and find the zeros?

OpenStudy (danjs):

yes solve for f'(x) = 0, those x values are the 'critical points' if any

OpenStudy (amenah8):

This is what i got when I factored the first D. Is it correct? 2(1 - x + 3x^2)=0

OpenStudy (danjs):

you would have to use the quadratic formula, any you will see there are no Real value answers, only complex, it is a parabola shifterd up 2, it never crosses x-axis (y=0)

OpenStudy (amenah8):

so i can say there are no intercepts? and no maximum or minimum?

OpenStudy (danjs):

so far nothing much, move to the second derivative and solve that one

OpenStudy (amenah8):

2nd D factored: 1(-1 + 6x) = 0 would this be the same situation as the 1st Derivative?

OpenStudy (danjs):

solve for the second derivative is zero 12x - 2 = 0

OpenStudy (amenah8):

so factor it, right? i took a 2 out of each and got: 2(6x - 1)

OpenStudy (danjs):

you get x=1/6, just solve it for x from algebra 1

OpenStudy (amenah8):

6x-1=0 --> add 1 to each side --> divide each side by 6 --> x=(1/6)

OpenStudy (danjs):

the second derivative tells you the rate of change of the first derivative, it is the concavity of the curve at any point if it exists

OpenStudy (amenah8):

so the inflection points, right?

OpenStudy (danjs):

it will tell you if the function is currently concave upwards or downwards at any point on the graph of f(x)

OpenStudy (danjs):

right

OpenStudy (amenah8):

so what does the (1/6) represent?

OpenStudy (danjs):

there could be one at that x=1/6 value, you have to test it in the second derivative, see if it changes sign there by using any value on either side of x=1/6 

OpenStudy (amenah8):

so plug (1/6) into the 2D equation?

OpenStudy (danjs):

x=1/6 , it represents where the second derivative is zero, inflection points, possible concavity change up/down down/up there on the graph

OpenStudy (danjs):

the two intervals to check are (-inf, 1/6) and (1/6 , +inf) any value in those will work to see the value of y'', negative - concave down positive - concave up i just picked x=0 and x=1, notice the sign of the second derivative changes at 1/6 from negative to positive

OpenStudy (amenah8):

where do you plug in the 0 and 1?

OpenStudy (danjs):

so from the two derivative tests you gained that the function has no max or min, it never changes increasing/decreasing second derivative told you there is an inflection point from concave down to concave up at x=1/6

OpenStudy (danjs):

into the second derivative, see if the value of that function changes sign +/- at x=1/6

OpenStudy (amenah8):

when plugging in 1: 10 when plugging in 0: -2

OpenStudy (danjs):

right, so the graph changes concave down(-) to concave up(+) at x=1/6

OpenStudy (amenah8):

and the 1st D told me there was no max or min, right? so now i just have to find the x and y intercepts to graph the original function

OpenStudy (danjs):

right, the first derivative never was zero, so the function is always increasing or always decreasing, never changes yeah i would find the axis intercept points

OpenStudy (amenah8):

do i use the original function for that? putting in 0 for x to find y-int and vice versa?

OpenStudy (danjs):

you see the first derivative must always be positive , a parabola vertex at + value and opens up, , always positive y values

OpenStudy (danjs):

that tells you the function is always increasing, since first derivative is always positive

OpenStudy (danjs):

overall info -axis intercepts points --inflection point y = f(1/6) = .. concave down to up -f(x) is always incrasing -domain of f(x) is any real number, range is also any real number

OpenStudy (amenah8):

so how do i find the x and y intercepts? is the vertex just (0,0), since the 1D told us there were no x intercepts?

OpenStudy (danjs):

y-intercept, when x=0 y=-1 , point (0,-1) x-intercept, when y=0 x=1/2, point (1/2, 0) inflection point, when x=1/6 y= -37/54 about -0.685

OpenStudy (amenah8):

did you plug those into the original equation?

OpenStudy (danjs):

if you plot those points, and know it is always increasing and changes concavity at that x=1/6, the general shape is